### v3

@@ Abstract
This paper presents a comparative analysis evaluating the accuracy of Large Language Models (LLMs) against traditional macro time series forecasting approaches. In recent times, LLMs have gained in popularity for forecasting due to their ability to capture intricate patterns in data and quickly adapt across very different domains. However, their effectiveness in forecasting macroeconomic time series data compared to conventional methods remains an area of interest. To address this, we conducted a rigorous evaluation of LLMs against traditional macro-prediction methods, using as common ground the FRED-MD database. Our findings provide valuable insights into the strengths and limitations of LLMs in the prediction of macroeconomic time series, shedding light on their applicability in real world scenarios.
@@ 1 Introduction
The recent emergence of Large Language Models (LLMs) has reshaped the landscape of natural language processing, ushering in a new era of computational linguistics. Bolstered by advancements in machine learning and deep neural networks, LLMs have garnered widespread attention for their remarkable ability to understand and generate human-like text. This transformative technology has revolutionized various applications, ranging from machine translation and sentiment analysis to chatbots and content generation. By leveraging vast amounts of text data and sophisticated algorithms, LLMs have demonstrated unparalleled proficiency in capturing linguistic nuances, contextual dependencies, and semantic meanings. [^1] For example, the abstract of this paper has been generated with a single prompt by chatGPT 3.5. As a result, they are increasingly becoming invaluable tools for researchers, developers, and businesses.
@@
Our first contribution in this paper is a thorough investigation on how TSLMs perform in predicting macroeconomic time series. A second contribution is a detailed comparison of their performance versus state-of-the-art time series methods such as Bayesian Vector Autoregressions (BVARs) and Factor Models as well as Bayesian Additive Regression Trees (BART), which has become a very popular nonparametric model for macroeconomic forecasting. In these endeavors, we focus on forecasting the variables contained in the FRED-MD dataset, a comprehensive repository of more than one hundred monthly macroeconomic variables curated by the Federal Reserve Economic Data (FRED) system. [^3] See https://research.stlouisfed.org/econ/mccracken/fred-databases/. This dataset encompasses a diverse array of economic indicators, and it is considered a benchmark in macroeconomic forecasting. A non-exhaustive lists of contributions using this data set to study the performance of alternative models in macroeconomic forecasting includes Stock and Watson (2006), Banbura et al. (2010), Carriero et al. (2016), Carriero et al. (2019), Korobilis and Pettenuzzo (2019), Koop et al. (2019), Chan (2022).
@@
As noted in Faria-e Castro and Leibovici (2024), there are some characteristics of LLM models that makes it problematic to use them in the pseudo out-of-sample forecasting exercise that is typically performed in the empirical macroeconomics literature. One major challenge is that LLMs are pretrained by developers on datasets of their choice, giving the researcher little to no control over the training data. For example, of the LLMs considered in this work, three out of five of them (including the best performing one, Moirai) list in their training data a large subset of the series we set out to forecast in this paper. To further complicate matters, since the training data lacks timestamps it’s hard to retrain the model up to a certain date, making it difficult to reproduce the type of real-time analysis that an economic forecaster is mostly concerned with. Solving these issues by retraining LLMs on subsets of data and adding data to the training set is usually not possible due to high hardware requirements. But having said that, it is important to recognize that even including the dataset of interest in real time would not remove the issue of a contaminated training set, i.e. including information that the forecaster would not have access to in real-time.
@@
Our empirical application is focused on forecasting the FRED monthly database using data ranging from 1960 to 2023. When working with traditional time series methods, we consider models of different cross-sectional dimensions, in order to understand the role that the information set has on the forecasting performance. The picture emerging from our analysis is one in which only two of the five TSLMs we consider are competitive against a simple AR benchmark (Salesforce’s Moirai and Google’s TimesFM). Moreover, when these more competitive TSLMs are stacked against Bayesian VARs and Factor Models, their forecasting performance is broadly comparable, if not slightly inferior. In particular, we find that the forecasting gains achieved by the econometric models tend to be more stable while TSLMs can perform very well for a handful of series but also show less reliability at times, as they seem to be more prone to generating the occasional unreasonable forecast. [^4] This result may not appear to some as totally surprising. When working with textual LLMs, we are sometimes exposed to “Model hallucinations”, i.e. the tendency of the model to generate content that is irrelevant, made-up, or inconsistent with the input data. On the other hand, TSLMs seem to work relatively better in the post-Covid-19 era (with the important caveat that the training set of these models does include information from the pandemic and post-pandemic period). While these results are based on using the TSLM models in a zero-shot forecasting manner, we also find that fine-tuning the TSLM models does not yield to significant improvements in their forecast accuracy. [^5] By “zero-shot”, we mean that the model is first trained on a large amount of data, and is then directly applied to solve a new task without re-training, or fine-tuning, its parameters on the new dataset.
@@
There are a few other papers that are looking at LLM in the macroeconomics and finance literature, and it is reasonable to expect that many more will follow. For example Bybee (2023) feed into an LLM model, OpenAI’s GPT-3.5, a historical sample of news articles from The Wall Street Journal (WSJ) and ask the model to predict various financial and macroeconomic quantities, which they then aggregate into a time-series of monthly and quarterly expectations that are compared to a variety of existing survey forecasts. Chen et al. (2022) feed global news text data from Thomson Reuters Real-time News Feed (RTRS) and Third Party Archive (3PTY) in 13 different languages into pretrained LLMs, namely BERT (developed by Google), RoBERTa (by Meta), and OPT (by Meta), and use the resulting sentiment scores to predict firm-level daily returns. Kim et al. (2024) feed standardized and anonymous financial statements to GPT4 and instruct the model to analyze them to determine the direction of future earnings. They find that LLMs outperforms financial analysts in their ability to predict earnings changes. These approaches focus on leveraging the typical LLMs abilities in natural language processing, while this paper is focused on the recent efforts of training LLMs specifically for time-series forecasting. Chen et al. (2025) study whether both DeepSeek and ChatGPT can be used to predict the stock market and find that a good news ratio extracted from prompting ChatGPT is positively correlated with contemporaneous market returns and significantly predicts subsequent returns up to six months out. The closest work we are aware of in the macroeconomics literature is Faria-e Castro and Leibovici (2024), which looks at time series LLMs, but only focuses on one LLM and one target variable (inflation). In contrast to this, in this paper we take on the ambition goal of looking at the performance of a variety of LLMs, comparing their accuracy against state of the art macroeconomic forecasting approaches, across a large set of macroeconomic indicators.
@@
The paper is organized as follows. Section 2 introduces LLMs and TSLMs and provides a discussion on how they are being leveraged to perform time series forecasting. Section 3 describes the econometric models we consider as state of the art benchmarks in our exercise. Section 4 describes the dataset and the design of the out-of-sample pseudo real time forecasting exercise and Section 5 presents our empirical results. Finally, Section 6 describes our fine-tuning process and the associated results, while Section 7 provides some concluding remarks and directions for further research. An Appendix provides technical details on the models considered and the data transformations used in the empirical analysis.
@@ 2.1 Time Series Language Models (TSLMs)
Inspired by these successes, researchers are now exploring the usefulness of LLMs in other settings. Along these lines, Lu et al. (2022) find that LLMs trained exclusively on textual data can effectively solve sequence modeling tasks in other modalities, such as numerical computation, vision, and protein fold prediction. Among these extensions, we are witnessing the emergence of Time Series Language Models (TSLM), bridging the gap between LLMs original text data training and the numerical nature of time series data. These models are sometimes referred to in the literature as Time Series Foundational Models (TSFM). TSLMs are now being used to accomplish a variety of time series related tasks, ranging from prediction and classification to anomaly detection and data imputation. Notable examples of this new wave of model development include Rasul et al. (2024), Goswami et al. (2024), Ekambaram et al. (2024), Garza and Mergenthaler-Canseco (2023), Das et al. (2024), and Ansari et al. (2024). At the core of this very recent advancements is a simple but fundamental point. While there are in principle significant differences between a textual LLM, whose main task is essentially to predict the next word, and a time series forecasting model, whose main objective is to predict the next value of one or multiple time series jointly, inherently these two tasks are deeply intertwined. They both aim at learning the sequential structure of the data from the past and present history of some features and with that attempt to predict a sequence of future outcomes. In the case of a TSLM, the pretraining is done on a possibly very large set of time series, $\bm{X}_{1:T}=\left(x_{1,1:T},...,x_{N,1:T}\right)$, and yields very large set of coefficients $\theta^{*}$ which describe a forecasting function $f_{\theta^{*}}$. Then, given $f_{\theta^{*}}$ and the current and past values of a time series of interest $y$ (not necessarily in the training set), the TSLM attempts to predict its future $h$ values $y_{T+1:T+h}$ using [^6] To keep the notation simple, we have assumed here a balanced set of features, i.e. all available over the same time period. If that is not the case, then we could instead write this as $\bm{X}_{1:T}=\left(x_{1,\underline{t}_{1}:T_{1}},...,x_{N,\underline{t}_{N}:T_{N}}\right)$, where $\underline{t}_{i}$ and $T_{i}$ would denote, respectively, the first and last observation of feature $i$, $i\in\left\{1,...,N\right\}$.
$\widehat{y}_{T+1:T+H|T}=f_{\theta^{*}}(\left.y_{1:T}\right)=f(\left.y_{1:T}|\bm{X}_{1:T},\theta^{*}\right).$ (1)
The main difference is that while natural language consists of words from a finite vocabulary, time series are real-valued.
@@ 2.2 Components of Time Series Language Models
The earliest applications of LLMs to time series forecasting were built on simply prompting existing textual LLMs with time series data. For example, PromptCast (Xue and Salim, 2023) relies on converting numerical time series into text prompts and forecasting in a sentence-to-sentence manner. [^7] Here is a weather forecasting example, to demonstrate how this approach would work: “From $t=1$ to $t=T$, the average temperature of region $i$ was $X_{i}$ degree on each day. What is the temperature going to be on day $T+h$?” There are multiple reasons why these approaches are sub-optimal. For once, these models requires a variable specific template to ask forecasting questions. In addition, using textual LLMs as is may lead to poor forecasting performances, mainly due to lack of real-world time series data in the training of these LLMs in the first place.
@@ 2.2.1 Tokenization
Patching is a process used to divide time series into shorter, fixed-length segments or windows called patches. These patches can be treated as tokens, capturing local patterns. For example, Moirai, TimesFM, TTM use patching to create tokens, while LagLlama augments each patch with lag features and time-based features to construct tokens. Importantly, the patch size is itself a hyper-parameter that can be adjusted based on the specific characteristics of the time series data. For example, larger patch sizes are more suitable for high-frequency time series, while smaller patch sizes tend to be more suitable for low frequency time series. Consider as an example a time series $x_{1:T}=\{4.7,4.76,6.8,7.2,6.1\}$. Tokenization with patching size$=3$ and overlap$=2$, will yield following tokens $\{4.7,4.76,6.8\}$, $\{4.76,6.8,7.2\}$, $\{6.8,7.2,6.1\}$. Without overlap the tokens will be $\{4.7,4.76,6.8\}$, $\{7.2,6.1\}$.
@@
Quantization is used to convert numerical time series values into a set of discrete tokens, similar to finite words in textual LLMs. This is achieved by dividing the value range of each time series, or a patch or set of patches (depending on the TSFM) into a predefined number of bins $\mathbb{B}$. Each data point is then assigned a token, a number between $\{1,\cdots\mathbb{B}\}$ based on the bin it falls into. There are different approaches to binning - uniform quantization divides the value range into bins of equal size, and data-dependent quantization adjusts the bin sizes based on the data distribution. Chronos (Ansari et al., 2024), for example, uses uniform quantization to encode time series into tokens. To continue with the previous example, let’s apply uniform quantization with $\mathbb{B}=4$ to the time series $x_{1:T}$. The mapping from $x_{1:T}$ to the four bins is carried out as follows:
@@
where $\tilde{x}_{t}$ is the scaled value. Different choices for $M$ and $S$ lead to the types of scaling used by the various TSFMs. For example, LagLlama uses $M=$ median of the context window and $S=$ inter-quartile range within the context window. [^10] A context window is the length of a time series considered as history for forecasting. It is also worth noting that scaling can be applied at the global level (using the value range of the entire time series), at a context window level, or at a patch level.

### v4

@@ Abstract
This paper presents a comparative analysis evaluating the accuracy of Time Series Language Models (TSLMs) against traditional macro time series forecasting approaches. In recent times, TSLMs have gained in popularity for forecasting due to their ability to capture intricate patterns in data and quickly adapt across very different domains. However, their effectiveness in forecasting macroeconomic time series data compared to conventional methods remains an area of interest. To address this, we conducted a rigorous evaluation of TSLMs against traditional macro-prediction methods, using as common ground the FRED-MD database. Our findings provide valuable insights into the strengths and limitations of TSLMs in the prediction of macroeconomic time series, shedding light on their applicability in real world scenarios. Building on these insights, we propose a novel hybrid approach that integrates the structure and interpretability of econometric models with the flexibility and nonlinearity of TSLMs, yielding the most robust and accurate forecasts across a range of horizons and variables.
@@ 1 Introduction
The recent emergence of Large Language Models (LLMs) has reshaped the landscape of natural language processing, ushering in a new era of computational linguistics. Bolstered by advancements in machine learning and deep neural networks, LLMs have garnered widespread attention for their remarkable ability to understand and generate human-like text. This transformative technology has revolutionized various applications, ranging from machine translation and sentiment analysis to chatbots and content generation. By leveraging vast amounts of text data and sophisticated algorithms, LLMs have demonstrated unparalleled proficiency in capturing linguistic nuances, contextual dependencies, and semantic meanings. [^1] For example, the first draft of the abstract of this paper was generated with a single prompt by chatGPT 3.5. As a result, they are increasingly becoming invaluable tools for researchers, developers, and businesses.
@@
Our first contribution in this paper is a thorough investigation on how TSLMs perform in predicting macroeconomic time series. A second contribution is a detailed comparison of their performance versus state-of-the-art time series methods such as Bayesian Vector Autoregressions (BVARs) and Factor Models, as well as Neural Networks and Bayesian Additive Regression Trees (BART), which have both become very popular nonlinear models for macroeconomic forecasting. A third contribution is the development of a novel hybrid approach that combines the structural rigor of traditional econometric models with the adaptability of TSLMs, leading to consistent improvements in forecast accuracy across variables and horizons. In these endeavors, we focus on forecasting the variables contained in the FRED-MD dataset, a comprehensive repository of more than one hundred monthly macroeconomic variables curated by the Federal Reserve Economic Data (FRED) system. [^3] See https://research.stlouisfed.org/econ/mccracken/fred-databases/. This dataset encompasses a diverse array of economic indicators, and it is considered a benchmark in macroeconomic forecasting. A non-exhaustive lists of contributions using this data set to study the performance of alternative models in macroeconomic forecasting includes Stock and Watson (2006), Banbura et al. (2010), Carriero et al. (2016), Carriero et al. (2019), Korobilis and Pettenuzzo (2019), Koop et al. (2019), Chan (2022).
@@
As noted in Faria-e Castro and Leibovici (2024), there are some characteristics of LLM models that makes it problematic to use them in the pseudo out-of-sample forecasting exercise that is typically performed in the empirical macroeconomics literature. One major challenge is that LLMs are pretrained by developers on datasets of their choice, giving the researcher little to no control over the training data. For example, of the LLMs considered in this work, three out of five of them (including the best performing one, Salesforce’s Moirai) list in their training data a large subset of the series we set out to forecast in this paper. To further complicate matters, since the training data lacks timestamps it’s hard to retrain the model up to a certain date, making it difficult to reproduce the type of real-time analysis that an economic forecaster is mostly concerned with. Solving these issues by retraining LLMs on subsets of data and adding data to the training set is usually not possible due to high hardware requirements. But having said that, it is important to recognize that even including the dataset of interest in real time would not remove the issue of a contaminated training set, i.e. including information that the forecaster would not have access to in real-time.
@@
More specifically, our empirical application is focused on forecasting all variables contained in the FRED monthly database using data ranging from 1960 to 2022. The picture emerging from our analysis is one in which only two of the five TSLMs we consider are competitive against a simple AR benchmark (Salesforce’s Moirai and Google’s TimesFM). Moreover, when these more competitive TSLMs are stacked against Bayesian VARs and Factor Models, their forecasting performance is broadly comparable, if not slightly inferior. In particular, we find that the forecasting gains achieved by the econometric models tend to be more stable while TSLMs can perform very well for a handful of series but also show less reliability at times, as they seem to be more prone to generating the occasional unreasonable forecast. [^4] This result may not appear to some as totally surprising. When working with textual LLMs, we are sometimes exposed to “Model hallucinations”, i.e. the tendency of the model to generate content that is irrelevant, made-up, or inconsistent with the input data. On the other hand, TSLMs seem to work relatively better in the post-Covid-19 era (with the important caveat that the training set of these models does include information from the pandemic and post-pandemic period). While these results are based on using the TSLM models in a zero-shot forecasting manner, we also find that fine-tuning the TSLM models does not yield to significant improvements in their forecast accuracy. [^5] By “zero-shot”, we mean that the model is first trained on a large amount of data, and is then directly applied to solve a new task without re-training, or fine-tuning, its parameters on the new dataset.
@@
There are a few other papers that are looking at LLM in the macroeconomics and finance literature, and it is reasonable to expect that many more will follow. For example Bybee (2023) feed into an LLM model, OpenAI’s GPT-3.5, a historical sample of news articles from The Wall Street Journal (WSJ) and ask the model to predict various financial and macroeconomic quantities, which they then aggregate into a time-series of monthly and quarterly expectations that are compared to a variety of existing survey forecasts. Chen et al. (2022) feed global news text data from Thomson Reuters Real-time News Feed (RTRS) and Third Party Archive (3PTY) in 13 different languages into pretrained LLMs, namely BERT (developed by Google), RoBERTa (by Meta), and OPT (by Meta), and use the resulting sentiment scores to predict firm-level daily returns. Kim et al. (2024) feed standardized and anonymous financial statements to GPT4 and instruct the model to analyze them to determine the direction of future earnings. They find that LLMs outperforms financial analysts in their ability to predict earnings changes. These approaches focus on leveraging the typical LLMs abilities in natural language processing, while this paper is focused on the recent efforts of training LLMs specifically for time-series forecasting. Chen et al. (2025) study whether both DeepSeek and ChatGPT can be used to predict the stock market and find that a good news ratio extracted from prompting ChatGPT is positively correlated with contemporaneous market returns and significantly predicts subsequent returns up to six months out. The closest work we are aware of in the macroeconomics literature is Faria-e Castro and Leibovici (2024), which looks at time series LLMs, but only focuses on one LLM and one target variable (inflation). In contrast to this, in this paper we take on the ambitious goal of looking at the performance of a variety of TSLMs, comparing their accuracy against state of the art macroeconomic forecasting approaches, across a large set of macroeconomic indicators.
@@
The paper is organized as follows. Section 2 introduces LLMs and TSLMs and provides a discussion on how they are being leveraged to perform time series forecasting. Section 3 describes the econometric models we consider as state of the art benchmarks in our exercise. Section 4 describes the dataset and the design of the out-of-sample pseudo real time forecasting exercise and Section 5 and Section 6 present our empirical results. Section 7 describes a novel hybrid approach that integrates econometric models with the flexibility and nonlinearity of TSLMs. Section 8 provides some concluding remarks and directions for further research. An Appendix provides technical details on the models considered, some additional results, and details on the data transformations used in the empirical analysis.
@@ 2.1 Time Series Language Models (TSLMs)
Inspired by these successes, researchers are now exploring the usefulness of LLMs in other settings. Along these lines, Lu et al. (2022) find that LLMs trained exclusively on textual data can effectively solve sequence modeling tasks in other modalities, such as numerical computation, vision, and protein fold prediction. Among these extensions, we are witnessing the emergence of Time Series Language Models (TSLM), bridging the gap between LLMs original text data training and the numerical nature of time series data. These models are sometimes referred to in the literature as Time Series Foundational Models (TSFM).
TSLMs are now being used to accomplish a variety of time series related tasks, ranging from prediction and classification to anomaly detection and data imputation. Notable examples of this new wave of model development include Rasul et al. (2024), Goswami et al. (2024), Ekambaram et al. (2024), Garza and Mergenthaler-Canseco (2023), Das et al. (2024), and Ansari et al. (2024). At the core of this very recent advancements is a simple but fundamental point. While there are in principle significant differences between a textual LLM, whose main task is essentially to predict the next word, and a time series forecasting model, whose main objective is to predict the next value of one or multiple time series jointly, inherently these two tasks are deeply intertwined. They both aim at learning the sequential structure of the data from the past and present history of some features and with that attempt to predict a sequence of future outcomes.
In the case of a TSLM, the pretraining is done on a possibly very large set of $N$ time series observed over the time periods $t=1,\dots,T$ which we collect in $\bm{X}_{1:T}=\left(x_{1,1:T},...,x_{N,1:T}\right)$. Pretraining yields a very large set of coefficients $\widehat{\theta}\in\Theta\subseteq\mathbb{R}^{d}$ which describe a forecasting function $f_{\widehat{\theta}}$. Then, given $f_{\widehat{\theta}}$ and the current and past values of a time series of interest $y_{1:T}$ (not necessarily in the training set), the TSLM attempts to predict its future $h$ values $y_{T+1:T+h}$ using [^6] To keep the notation simple, we have assumed here a balanced set of features, i.e. all available over the same time period. If that is not the case, then we could instead write this as $\bm{X}_{1:T}=\left(x_{1,\underline{t}_{1}:T_{1}},...,x_{N,\underline{t}_{N}:T_{N}}\right)$, where $\underline{t}_{i}$ and $T_{i}$ would denote, respectively, the first and last observation of feature $i$, $i\in\left\{1,...,N\right\}$.
[…]
The main difference with LLMs is that while natural language consists of words from a finite vocabulary, time series are real-valued.
@@ 2.2 Components of Time Series Language Models
The earliest applications of LLMs to time series forecasting were built on simply prompting existing textual LLMs with time series data. For example, PromptCast (Xue and Salim, 2023) relies on converting numerical time series into text prompts and forecasting in a sentence-to-sentence manner. [^7] Here is a weather forecasting example, to demonstrate how this approach would work: “From $t=1$ to $t=T$, the average temperature of region $i$ was $X_{i}$ degree on each day. What is the temperature going to be on day T+hT+h?” There are multiple reasons why these approaches are sub-optimal. For once, these models requires a variable specific template to ask forecasting questions. In addition, using textual LLMs as is may lead to poor forecasting performances, mainly due to lack of real-world time series data in the training of these LLMs in the first place.
@@ 2.2.1 Tokenization
Patching is a process used to divide time series into shorter, fixed-length segments or windows called patches. These patches can be treated as tokens, capturing local patterns. For example, Moirai, TimesFM, TTM use patching to create tokens, while LagLlama augments each patch with lag features and time-based features to construct tokens. Importantly, the patch size is itself a hyper-parameter that can be adjusted based on the specific characteristics of the time series data. For example, larger patch sizes are more suitable for high-frequency time series, while smaller patch sizes tend to be more suitable for low frequency time series. Consider as an example a time series $x_{1:T}=\{4.7,4.76,6.8,7.2,6.1\}$. Tokenization with patching size =3=3 and overlap =2=2, will yield following tokens $\{4.7,4.76,6.8\}$, $\{4.76,6.8,7.2\}$, $\{6.8,7.2,6.1\}$. Without overlap the tokens will be $\{4.7,4.76,6.8\}$, $\{7.2,6.1\}$.
@@
Quantization is used to convert numerical time series values into a set of discrete tokens, similar to finite words in textual LLMs. This is achieved by dividing the value range of each time series, or a patch or set of patches (depending on the TSLM) into a predefined number of bins $\mathbb{B}$. Each data point is then assigned a token, a number between $\{1,\cdots\mathbb{B}\}$ based on the bin it falls into. There are different approaches to binning - uniform quantization divides the value range into bins of equal size, and data-dependent quantization adjusts the bin sizes based on the data distribution. Chronos (Ansari et al., 2024), for example, uses uniform quantization to encode time series into tokens. To continue with the previous example, let’s apply uniform quantization with $\mathbb{B}=4$ to the time series $x_{1:T}$. The mapping from $x_{1:T}$ to the four bins is carried out as follows:
@@
where $\tilde{x}_{t}$ is the scaled value. Different choices for $M$ and $S$ lead to the types of scaling used by the various TSLMs. For example, LagLlama uses $M=$ median of the context window and $S=$ inter-quartile range within the context window. [^10] A context window is the length of a time series considered as history for forecasting. It is also worth noting that scaling can be applied at the global level (using the value range of the entire time series), at a context window level, or at a patch level.
Embeddings
In TSLMs, tokenization — through scaling, patching, or quantization — converts raw sequence data into discrete tokens. This process is similar to how textual LLMs tokenize sentences into words or subwords. The time series tokens are then transformed into embeddings, which are dense numerical vectors. Textual embeddings capture semantic relationships between words, while in TSLM embeddings encode local temporal patterns such as trend or seasonality. This transformation occurs via an embedding layer, a learnable component of the transformer architecture (we discuss the architecture in the next subsection), which maps tokens to a vector space suitable for processing by attention mechanisms. While architecturally part of the transformer, the embedding process creates inputs capturing the contextual information of the time series data for the transformers.